Circular No. 6976

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1998cx IN NGC 6209 A. Piemonte, Pontificia Universidad Catolica; and F. Patat, European

Southern Observatory (ESO), report that they obtained a CCD spectrogram (60-min exposure, range 315–970 nm, resolution 1.5 nm) of SN 1998cx (cf. IAUC 6970) with the ESO 1.5-m telescope (+ Boller & Chivens spectrograph) on July 22.08 UT at La Silla. A full reduction of the spectrum shows that this supernova is typical of type-Ia events at intermediate phases; it is dominated by Fe II lines, although some features arising from intermediate elements (Ca II and Si II) are still present. This and the expansion velocity deduced from Ca H and K and the Ca II infrared triplet (10 800 km/s) suggest that SN 1998cx has been observed around 5–6 weeks after maximum light. The spectrum is very similar to that shown by SN 1994D at similar phases (Patat et al. 1995, MNRAS 278, 111). The recession velocity of the parent galaxy is 5865 km/s (Mathewson et al. 1992, Ap.J. Suppl. 81, 413).

XTE J2123–058 R. I. Hynes, University of Sussex; P. A. Charles, University of Oxford;

C. A. Haswell, University of Sussex; and J. Casares, M. Serra-Ricart, and C. Zurita, Instituto de Astrofisica de Canarias, report: “Time-resolved spectroscopy (range 350–685 nm) of the x-ray transient XTE J2123–058 was obtained with the 4.2-m William Herschel Telescope at La Palma on July 19–20. The spectra reveal a nearly featureless continuum with only C III/N III 464.0-nm (average equivalent width 0.20 nm) and He II 468.6-nm (equivalent width 0.23 nm) emission being prominent. He II 468.6-nm shows a radial-velocity modulation with a period of 6.05 ± 0.10 hr, consistent with the photometric periods of Tomsick et al. (IAUC 6972) and Ilovaisky and Chevalier (IAUC 6975) and confirming that the photometric period is indeed the orbital period of the system. The velocity semi-amplitude is ∼ 180 km/s, with maximum radial velocity close to photometric phase 0.75, suggesting an association with the compact object or disk. Further monitoring of this system is encouraged.”

NOVA OPHIUCHI 1998 Additional CCD photometry by D. Hanžl, Brno, Czech Republic (cf.

IAUC 6943, 6955): June 29.891 UT, V = 13.57 ± 0.06, B − V = +0.68 ± 0.1; July 3.905, V = 13.85 ± 0.03, –; 18.874, V = 14.41 ± 0.06, –; 20.886, V = 14.76 ± 0.06, –.

1998 July 23  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green